# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v5

@@ 5 Related research and a brief history
We recently became aware of an independent discovery of the 8-conic theorem [Eid13], Sect. 9.2. (English translation here.) The author Morten Eide has in the meantime become a co-author of this article and contributed Sec. 4.
@@
The impetus for this article came from a podcast featuring Roger Penrose [Har20], in which he describes formulating the 8-conic theorem in the early 1950’s, but not having found the time to publish it.
@@
In 2024 the authors of [CGN] contacted Roger Penrose to report their proof of the theorem. A collaboration ensued, in which Roger Penrose shared details of his original approaches to the proof of the theorem. He recalled that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Later, as a graduate student, he discovered an approach to prove the 8-conic theorem using an analogous 8-quadric theorem in 3-D. It is this approach that is reconstructed in the proof in Sec. 3.
Related research has produced two companion articles that provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5-D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9-D space of quadrics. The proofs have a similar structure to each other; the advantages of the 3-D parameter space approach over the 2-D approach become clear in the final stages of the proofs.
6 Outlook for future work
@@
The possibility of gluing several Penrose cubes together, described above, remains an interesting option to obtain further theorems. For example, this method can be used to extend Poncelet’s theorem for triangles to general polygons. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.
Regarding cube configurations with degenerate conics or quadrics, further conditions are needed to ensure the validity of the main theorems presented in this article. This can be investigated in parameter space by looking at limits of regular configurations. It seems very worthwhile to us to further develop these techniques making use of the projective geometric structures of parameter space, as done in [CGN] and [ACG+].

### v7

@@ 8 Related research and a brief history
We recently became aware of an independent discovery of the 8-conic theorem [Eid13], Sect. 9.2. (English translation here.) The author, Morten Eide, has in the meantime become a co-author of this article.
@@
The seed for this article was planted when one of the authors heard a podcast featuring Roger Penrose [Har20], in which he describes formulating the 8-conic theorem in the early 1950’s, but not having found the time to publish it.
@@
In May, 2024 some of the current authors contacted Roger Penrose to report that they had obtained a proof of the theorem. A collaboration ensued in which Roger Penrose shared details of his original approaches to the proof of the theorem. He recalled that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Later, as a graduate student, he discovered an approach to prove the 8-conic theorem using an analogous 8-quadric theorem in 3-D. It is this approach that is reconstructed in the geometric proof in Sec. 4. The algebraic proof presented in Sec. 6 is due to Morton Eide.
Our initial proof of the 8-conic theorem, in contrast, worked in the 5-D parameter space of conics. After the proof of the 8-quadric theorem in 3-D (presented in this article) was obtained, we also developed a proof in the 9-D parameter space of quadrics. The parameter-space proofs have a similar structure to each other, but just as in the proofs presented in the first part of this article, the parameter-space proof of the 8-quadric theorem is shorter and easier than the parameter-space proof of the 8-conic theorem. We hope to present these proofs in the near future in one or two further preprints.
9 Recapitulation and Outlook
Beginning with a series of examples of familiar theorems in projective geometry we formulated the Penrose 8-conic theorem, that includes these examples as special cases, and proved it, first geometrically and then algebraically. Here we attempt to evaluate the significance of the theorem and indicate some directions for future research. Much but not all of the following discussion can be transferred also to the associated Penrose 8-quadric theorem.
We saw that the theorem has a striking universality, so that many theorems of projective geometry can be positioned as special cases. Furthermore, the determinant formulas in Section 7 provide an explicit parametrization of all valid Penrose cubes, yielding a (17-dimensional) space whose points each represent a Penrose cube. Since such a cube can be interpreted as a theorem, any path between two points in the space represents a continuous metamorphosis of one theorem to the the other, in which all intermediate stages are also theorems. This is a certainly a remarkable space.
One general question is: Which projective theorems correspond to points in this Penrose space and which not? That is, how much of projective geometry can be seen as a consequence of the Penrose 8-conic theorem? How many different “theorems” does it contain? To answer this, observe that it is divided into regions by sub-manifolds where the rank of a matrix of one of the 8 conics changes. Each region can be considered to represent one equivalence class of theorems. Factoring out by different labeling of the same cube can be used to further reduce the number of equivalence classes. The task remains daunting.
Our approach has been more “bottom-up” by beginning with specific theorems and searching for patterns there. For example, the Penrose cubes for many of known theorems have a 3-fold symmetry, whereby the 4 levels of a Penrose cube (balanced on its bottom vertex $\mathbf{{T}}^{0}$) each have the same rank and type (point-wise or line-wise). Consult Figure 12. We have carried out a survey of 13 or 14 possible types, producing the beginning of a classification that will be published separately.
Besides the rank of the conics, other factors that produce variation in the specialization include:
Distinction of point-wise and line-wise conics
The conics of the Penrose cube are complete conics (Section 3.3.2), including both a point-wise and line-wise partner. Note that dualizing a Penrose theorem is obtained by dualizing its Penrose cube. The formulas of Section 6 necessarily apply to either point-wise or line-wise conics. The interaction between these two aspects can be subtle and deserves further attention, especially the rare cases that one or the other of the partners vanishes.
Real and complex elements
Our theory considers conics in $\mathbb{R}P^{2}$, whose polynomials have real coefficients. But the zero sets of these polynomials can become imaginary in manifold ways. Tracking this transition is an important aspect that deserves further attention. The presence of imaginary conics also allows connections to euclidean geometry, as the next item explains.
Interaction with metric geometry
The Cayley-Klein construction of metric spaces within projective space can find expression in a Penrose cube since the Cayley-Klein metric absolute is a conic. For example, a euclidean circle is a conic that has double contact with the euclidean absolute imaginary points $(I,J)$. Including this point pair in a Penrose cube guarantees that its neighbors (when regular) will be circles. The Monge Circle theorem discussed above is one of many examples. On the other hand, euclidean confocal conics are line-wise conics with a particular relationship to the euclidean absolute that allows Penrose cubes to be contructed expressing Euclidean theorems about confocal conics and focal points. Such examples raise the question to what extent the domain of the Penrose 8-conic theorem extends also into metric geometry.
@@ 9 Recapitulation and Outlook
We have also explored extensions of the 8-conic theorem. The possibility of gluing several Penrose cubes together, mentioned above, remains an interesting option to obtain further theorems. For example, this method can be used to construct Pocelet configurations for general polygons out of those for triangles. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.
We also saw that the determinant theory of Section 7 can be directly extended to other configurations, such as hyper-cubes. There are also outstanding questions regarding the finer points of the subdeterminants of $\mathbf{{S}}_{\{1,2,3\}}$; for example, to work out in detail how the simultaneous vanishing of several subdeterminants can effect in subtle ways the dual partners of particular conics in the cube. Another, more general direction is, in the underlying matrices, to replace the linear forms $\mathbf{p}_{i}$ with $n-$order curves; conics will be replaced with $2n$-order curves, whereby the number of contact points is $2n^{2}$.
A more general direction of research involves the place of the Penrose 8-conic theorem within projective geometry. It is an archetypal theorem, containing within itself innumerable known and not-yet-known theorems. Based on our experience in the process of preparing this article, it has the potential to enrich how projective geometry is practiced and perceived, expanding the focus from isolated theorems to a continuum of metamorphoses of which the theorems themselves are static snapshots.
Appendix A Appendix
In this appendix, we show the existence of two quadrics completing the cube in the case where all six ring planes concur in a common axis. By the results of sections 6 and 7, only one of the two will be attainable as the limit of cubes arising in the general case, where the ring planes don’t have a common axis (see remark below).
In the case under consideration, when all ring planes meet in a common axis $\Omega$, all of the quadrics will be in double contact, i.e., share two contact elements: $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{2}$ share the points
$\mathbf{{R}}_{0}^{1}\cap\mathbf{{R}}_{0}^{2}=\mathbf{{R}}_{0}^{1}\cap\Omega$ (93)
and, by marching around the cube in the same fashion as Lemma 12, these two points are common to all the quadrics of the cube:
$\mathbf{{R}}_{i}^{j}\cap\Omega=\mathbf{{R}}_{k}^{m}\cap\Omega,$ (94)
for all indices $i,j,k,m$, as are the tangent planes at these points (the case where $\Omega$ is tangent to the rings provides some interesting variations but we omit consideration of it as it doesn’t differ in the essentials).
In an appropriate basis, the matrices of the family of quadrics sharing these two contact elements can be written
$\mathbf{{Q}}=a_{0}\mathbf{{Q}}_{0}+a_{1}\mathbf{{Q}}_{1}+a_{2}\mathbf{{Q}}_{2}+a_{3}\mathbf{{Q}}_{3}$ (95)
where
$\mathbf{{Q}}_{0}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ \end{pmatrix},\qquad\mathbf{{Q}}_{1}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix},$
$\mathbf{{Q}}_{2}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\\ \end{pmatrix},\qquad\mathbf{{Q}}_{3}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ \end{pmatrix}.$
In order for the contact elements to be contained in the literal sense, we must have $a_{0}=0$, thus this parametric space can be viewed as an affine $3$-space with the quadrics $\mathbf{{Q}}$ given by equation (95) as ‘points’. We shall use single quotes when talking about geometric objects in this space in order to distinguish them from geometric elements in the space in which the $\mathbf{{Q}}$ appear as surfaces, e.g., we shall refer to a pencil of quadrics lying in this parametric space as a ‘line’. It makes sense to include also the ‘plane’ $a_{0}=0$, making the parametric space into a projective space. This ‘plane’ consists of the plane pairs which meet in $\Omega$. The double planes form a ‘conic’ $\mathtt{V}$, the locus of ‘points’ $\mathbf{{Q}}$ satisfying
$a_{0}=0,\qquad a_{1}a_{2}-a_{3}^{2}=0.$ (96)
In summary, when all the ring planes share an axis, all the quadrics of the cube can be viewed as belonging to a parametric $3$-space, including any possible choice of $\mathbf{{T}}^{0}$ that completes the cube. Any quadric in this space and in ring contact with a given quadric $\mathbf{{Q}}$ lies on a ‘line’ containing a ‘point’ of $\mathtt{V}$. Therefore, the family of quadrics in ring contact with $\mathbf{{Q}}$ is the ‘cone’ $\mathtt{C}_{\mathbf{{Q}}}$ made up of the ‘lines’ connecting $\mathbf{{Q}}$ to any ‘point’ of $\mathtt{V}$. Any possible $\mathbf{{T}}^{0}$ must therefore be in the intersection
$\mathtt{C}_{\mathbf{{S}}_{1}}\cap\mathtt{C}_{\mathbf{{S}}_{2}}\cap\mathtt{C}_{\mathbf{{S}}_{3}}.$ (97)
The intersection of two ‘cones’ in general forms a fourth degree curves, however, in the present case, they meet in the ‘conic’ $\mathtt{V}$ and thus there intersection breaks down into a pair of ‘conics’:
$\mathtt{C}_{\mathbf{{S}}_{i}}\cap\mathtt{C}_{\mathbf{{S}}_{j}}=\mathtt{V}\cup\mathtt{W}_{ij}.$ (98)
Therefore
$\mathtt{C}_{\mathbf{{S}}_{1}}\cap\mathtt{C}_{\mathbf{{S}}_{2}}\cap\mathtt{C}_{\mathbf{{S}}_{3}}=\mathtt{V}\cup(\mathtt{W}_{12}\cap\mathtt{W}_{13}).$ (99)
In general, $\mathtt{W}_{12}\cap\mathtt{W}_{13}$ consists of two ‘points’ either of which can be taken as a regular quadric completing the cube. Figure 19 illustrates this for the case of conics, where the condition of the planes sharing a common axis is replaced by that of the face points of all faces coinciding (see [CGN25] for more on these parametric three spaces in the conic case).
Figure 19: Penrose configuration where all face points coincide and, as a result, the cube has two completions. $\mathbf{{S}}_{0}$ is orange, $\mathbf{{T}}^{1}$, $\mathbf{{T}}^{2}$, $\mathbf{{T}}^{3}$ are blue, $\mathbf{{S}}_{1}$, $\mathbf{{S}}_{2}$, $\mathbf{{S}}_{3}$ are magenta, and the two possible completions $\mathbf{{T}}^{0}$ are dark green and lime green.
Remark 21.
The approach of sections 6 and 7 always provides a unique completion regardless of whether or not the face axes/face points collapse to one. Therefore the two completions provided in the present appendix can be distinguished: one that arises from the determinant formulas and is thus stable under parametric perturbations which disturb the coincident face-axes/face-point property, and one which is only present thanks to this coincidence. In figure 19 the former is the dark green central circle and the latter is the lime green circle surrounding the magenta conics.